Financial data for Joel de Paris, Inc., for last year follow:

Joel de Paris, Inc.
Balance Sheet
Beginning
Balance
Ending
Balance
Assets
Cash $ 128,000 $ 127,000
Accounts receivable 333,000 476,000
Inventory 563,000 479,000
Plant and equipment, net 816,000 798,000
Investment in Buisson, S.A. 395,000 429,000
Land (undeveloped) 252,000 245,000
Total assets $ 2,487,000 $ 2,554,000
Liabilities and Stockholders' Equity
Accounts payable $ 373,000 $ 339,000
Long-term debt 1,016,000 1,016,000
Stockholders' equity 1,098,000 1,199,000
Total liabilities and stockholders' equity $ 2,487,000 $ 2,554,000


Joel de Paris, Inc.
Income Statement
Sales $ 4,650,000
Operating expenses 3,859,500
Net operating income 790,500
Interest and taxes:
Interest expense $ 125,000
Tax expense 193,000 318,000
Net income $ 472,500


The company paid dividends of $371,500 last year. The “Investment in Buisson, S.A.,” on the balance sheet represents an investment in the stock of another company. The company's minimum required rate of return of 15%.

Required:

1. Compute the company's average operating assets for last year.

2. Compute the company’s margin, turnover, and return on investment (ROI) for last year. (Round "Margin", "Turnover" and "ROI" to 2 decimal places.)

3. What was the company’s residual income last year?
Answer #1

Average operating assets

= [(2,487,000-252,000-395,000)+(2,554,000-245,000-429,000)]/2

= 1,860,000

Margin = Net operating income/sales

= 790,500/4,650,000

= 17%

.

Turnover = Sales/Average operating assets

= 4,650,000/1,860,000

= 2.5

.

ROI = Margin*Turnover

= 17*2.5

= 42.50%

Residual income = Net operating income - required return

= 790,500-(1,860,000*15%)

= 511,500

Comment if you face any issues